# Sparse Hanson-Wright Inequalities with Applications

### v2

@@ 4 Proof of Theorem 1
Let $r$ be an even integer. Let $X_{1},\cdots,X_{n}$ be independent centered sparse sub-gaussian random variables, i.e. $X_{i}=x_{i}\xi_{i}$, where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$ with a constant $K>0$. Let $A$ be a diagonal-free $n\times n$ matrix. Then
$\mathbb{E}\left|\sum_{i\neq j}a_{ij}X_{i}X_{j}\right|^{r}\leq(CK^{2})^{r}\max\left\{\left(r\sum_{i,j}a_{ij}^{2}p_{i}p_{j}\right)^{r/2},{(r\gamma_{1,\infty})^{r}},{(r^{2}\|A\|_{\max})^{r}}\right\}.$ (4.1)
With Lemma 4.1, we now proceed to prove our main result.
Proof of Theorem 1.
$\leq\left(\frac{1}{t}\max\left\{CK^{2}\sqrt{r\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},CK^{2}r\gamma_{1,\infty},CK^{2}r^{2}\|A\|_{\max}\right\}\right)^{r}.$ (4.3)
For any $t>0$, if
$\min\left\{\frac{t^{2}}{Ce^{2}K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{CeK^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{CeK^{2}\|A\|_{\max}}}\right\}<2,$ (4.4)
then in the inequality in Theorem 1 there is
$\exp\left(-\min\left\{\frac{t^{2}}{Ce^{2}K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{CeK^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{CeK^{2}\|A\|_{\max}}}\right\}\right)\geq e^{-2}.$
We can choose constant $C^{\prime}=e^{2}$ such that the right-hand side in Theorem 1 is at least 1 and the theorem holds.
$2\leq r\leq\min\left\{\frac{t^{2}}{Ce^{2}K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{CeK^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{CeK^{2}\|A\|_{\max}}}\right\},$
we have that
$\frac{1}{t}\max\left\{CK^{2}\sqrt{r\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},CK^{2}r\gamma_{1,\infty},CK^{2}r^{2}\|A\|_{\max}\right\}\leq\frac{1}{e},$
and further
$\mathbb{P}\left(\left|\sum_{i\not=j}a_{ij}X_{i}X_{j}\right|\geq t\right)$ $\leq e^{-r}$
$\leq\exp\left(-\left(c\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{K^{2}\|A\|_{\max}}}\right\}-2\right)\right)$
$=e^{2}\exp\left(-c\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{K^{2}\|A\|_{\max}}}\right\}\right).$
4.1 Proof of Lemma 4.1
In the remainder of this paper, we use $C$ to represent an absolute constant whose value may change from line to line. We do not keep track of these absolute constants explicitly.
@@
    A multigraph $G(\Pi)$ is obtained by identifying the vertices in $\mathcal{G}$ according to the partition $\Pi$. The multigraph $G(\Pi)$ has $r$ edges and $k$ vertices, and each vertex has a degree of at least 2. Also, we focus on the ones that are loop-free, since $A$ is diagonal-free and any loop in $G(\Pi)$ gives a zero in the sum.
@@
$\mathbb{E}|x_{i}|^{l}=p_{i}\quad\text{for }l\geq 2,$
and
$\mathbb{E}|X_{i}|^{l}=\mathbb{E}|x_{i}|^{l}\mathbb{E}|\xi_{i}|^{l}\leq p_{i}(K{l})^{l}\leq p_{i}(eK)^{l}l!,$
@@
The proof of Lemma 4.2 is deferred to the end of this section. It is worth noting that the bound on $\mathcal{S}_{G(\Pi)}$ in Lemma 4.2 is uniform and independent of $G(\Pi)$. To proceed from (4.8), we now need to estimate
@@
    $\sum_{\phi_{i}:[2r_{i}]\to[k_{i}]}\prod_{\tau=1}^{k}\deg_{G(\phi_{i})}(\tau)!$ $\leq\sum_{d_{1},\dots,d_{k_{i}}}\binom{2r_{i}}{d_{1}+\dots+d_{k_{i}}=2r_{i}}\cdot\prod_{\ell=1}^{k_{i}}d_{\ell}!$ (4.14)
@@
    $\leq 2^{4r_{i}}r_{i}^{2r_{i}}.$ (4.16)
Therefore, combining the steps above, with $r_{i}\geq 2$ we have
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ $\leq\frac{2^{r+k}}{k!}\max_{\text{component partition}}\left(\prod_{i=1}^{t}\binom{r}{r_{i}-1}\binom{k}{k_{i}}\right)\left(\prod_{i=1}^{t}2^{4r_{i}}r_{i}^{2r_{i}}\right)$
$\leq\frac{2^{r+k}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{er}{r_{i}-1}\right)^{r_{i}-1}\left(\frac{ek}{k_{i}}\right)^{k_{i}}2^{4r_{i}}r_{i}^{2r_{i}}$
$\leq\frac{C^{r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r}{r_{i}}\right)^{r_{i}-1}\left(\frac{k}{k_{i}}\right)^{k_{i}}r_{i}^{2r_{i}}$
$\leq\frac{C^{r}r^{2r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r}{r_{i}}\right)^{r_{i}-1}\left(\frac{k}{k_{i}}\right)^{k_{i}}\left(\frac{r_{i}}{r}\right)^{2r_{i}}$
$\underset{(a)}{\leq}\frac{C^{r}r^{2r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r_{i}}{r}\right)^{r_{i}-k_{i}+1}.$ (4.17)
@@
Here, the last inequality follows from Lemma A.2 and that the maximum of $\prod_{i}z_{i}^{z_{i}}$ is attained when $t-1$ of the $z_{i}$’s are equal to 1, and the remaining $z_{i}$ equals $Z-(t-1)$. Therefore, we have
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ $\leq\frac{C^{r}r^{2r}}{k!}\cdot\frac{(r-k+1)^{r-k+1}}{(r-k+t)^{r-k+t}}$ (4.19)
$\leq C^{r}r^{2r}k^{-k}{t^{-(t-1)}}$ (4.20)
$\leq C^{r}r^{2r-k-t+1}\left(\frac{r}{k}\right)^{k}\left(\frac{r}{t}\right)^{t}$ (4.21)
$\leq C^{r}r^{2r-k-t}.$ (4.22)
@@
$\leq(CK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}\cdot r^{2r-k-t}$ (4.25)
$\leq C^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}(r^{2}K^{2}\|A\|_{\max})^{r-k}(rK^{2}\gamma_{1,\infty})^{k-2t}(\sqrt{r}K^{2}\sigma)^{2t}$ (4.26)
$\leq C^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\max\left\{\frac{(r^{2}K^{2}\|A\|_{\max})^{r}}{r/(r-k)},\frac{(rK^{2}\gamma_{1,\infty})^{r}}{r/(k-2t)},\frac{(\sqrt{r}K^{2}\sigma)^{r}}{r/2t}\right\}$ (4.27)
$\leq C^{2r}r^{2}\max\left\{{(r^{2}K^{2}\|A\|_{\max})^{r}},{(rK^{2}\gamma_{1,\infty})^{r}},{(\sqrt{r}K^{2}\sigma)^{r}}\right\}$ (4.28)
$\leq C_{1}^{r}\max\left\{{(r^{2}K^{2}\|A\|_{\max})^{r}},{(rK^{2}\gamma_{1,\infty})^{r}},{(\sqrt{r}K^{2}\sigma)^{r}}\right\},$ (4.29)
where $C,C_{1}$ are absolute constants.
4.2 Proof of Lemma 4.2
@@
$\mathcal{S}_{G}:=\sum_{s\in[n]^{V(G)}}\left(\prod_{\{\pi_{1},\pi_{2}\}\in E(G)}a_{s_{\pi_{1}}s_{\pi_{2}}}\cdot\prod_{\pi\in V(G)}p_{s_{\pi}}\right).$ (4.30)
@@
    For each deleted edge, say $\{u,v\}$, we bound the corresponding $a_{uv}\leq\|A\|_{\max}$, and we will delete $r-k$ many edges this way in total, which implies that we delete a total of $r-k$ edges from $G$. These deleted edges collectively contribute a factor of $\|A\|_{\max}^{r-k}$ to bound $\mathcal{S}_{G}$, i.e.
@@
4.3 Proof of Theorem 2
The proof of Theorem 2 follows a similar moment estimation scheme to that of Theorem 1. We provide a brief proof here, focusing on the differences. First of all, the moments of $\alpha$-sub-exponential random variables satisfy
$\mathbb{E}|\xi_{i}|^{l}\leq K^{l}l^{\frac{l}{\alpha}}.$
For $\alpha\geq 1$, we simply bound $\mathbb{E}|\xi_{i}|^{l}\leq K^{l}l^{l}$, which leads to results identical to those of Theorem 1. We will focus on the case where $\alpha\in(0,1)$. Since
$\mathbb{E}|X_{i}|^{l}=\mathbb{E}|x_{i}|^{l}\mathbb{E}|\xi_{i}|^{l}\leq p_{i}(eK)^{l}l!\cdot l^{(\frac{1}{\alpha}-1)l},$
@@ 4.3 Proof of Theorem 2
$\mathbb{E}|X^{T}AX|^{r}$ $\leq(CK^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}r^{2r-k-t}\cdot\big{(}2^{2r}r^{2(r-k+1)}\big{)}^{\frac{1}{\alpha}-1}$ (4.40)
$\leq(C_{\alpha}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}r^{2r-k-t}\cdot r^{2(\frac{1}{\alpha}-1)(r-k+1)}$ (4.41)
$\leq(C_{\alpha}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}\cdot r^{\frac{2}{\alpha}r-(\frac{2}{\alpha}-1)k-t}$ (4.42)
$=(C_{\alpha}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}(r^{\frac{2}{\alpha}}\|A\|_{\max})^{r-k}(r\gamma_{1,\infty})^{k-2t}(\sqrt{r}\sigma)^{2t}$ (4.43)
$\leq(C_{\alpha}K^{2})^{r}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}^{r}.$ (4.44)
Here in the last inequality, we use Young’s inequality for the products of three terms.
Hence,
@@
for any even integer $r$.
From the above norm bound, by similar discussion in the proof to Theorem 1, we derive the tail inequality. By Markov’s inequality,
$\mathbb{P}\left(\left|\sum_{i\not=j}a_{ij}X_{i}X_{j}\right|\geq t\right)\leq\left(\frac{1}{t}C_{\alpha}K^{2}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}\right)^{r}.$ (4.45)
We choose $r$ to be the largest even integer such that
$2\leq r\leq\min\left\{\frac{t^{2}}{C_{\alpha}e^{2}K^{4}\sigma^{2}},\frac{t}{C_{\alpha}eK^{2}\gamma_{1,\infty}},\left(\frac{t}{C_{\alpha}eK^{2}\|A\|_{\max}}\right)^{\frac{\alpha}{2}}\right\}$ (4.46)
and this implies
$\frac{1}{t}C_{\alpha}K^{2}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}\leq\frac{1}{e}.$ (4.47)
If the right-hand side of (4.46) is less than 2, then we choose $C^{\prime}=e^{2}$ such that the conclusion of Theorem 2 trivially holds. This proves Theorem 2.
5 Proofs of Theorem 3 and Corollary 1
Following the proof of the off-diagonal case, we now focus on the concentration of the linear form. Note that given Theorem 3, Corollary 1 immediately holds as
$\mathbb{E}|X_{i}^{2}-\mathbb{E}X_{i}^{2}|^{l}\leq 2^{l}\mathbb{E}|X_{i}|^{2l}\leq 2^{l}p_{i}K^{2l}(2l)^{\frac{2l}{\alpha}}.$ (5.1)
In the rest of this section, we will prove Theorem 3 by moment method. The following lemma provide a moment bound for the linear sum.

### v3

@@ 5 Proofs of Theorem 1 and Theorem 2
Under the assumptions of Theorem 1, let $r$ be any even integer. We have
$\|X^{T}AX\|_{L_{r}}\leq C_{\alpha}K^{2}\max\left\{r^{\max\{2,\frac{2}{\alpha}\}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\},$ (5.1)
where $\gamma_{1,\infty}$ is defined in (4.1) and $\sigma:=\sqrt{\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}}.$ Here, $C_{\alpha}=48e^{5+2/e}\cdot 8^{\max\{1,1/\alpha\}}$.
To enhance the clarity of our presentation, we structure our proofs as follows. In Section 5.1, we provide the proof of Lemma 5.1 for the special case where $\alpha=1$. Subsequently, in Section 5.2, we outline the necessary modifications to extend the conclusion to the general case of $\alpha\in(0,\infty]$. Finally, in Section 5.3, we demonstrate that Theorem 1 follows as a direct consequence of Lemma 5.1.
5.1 Proof of Lemma 5.1 when $\alpha=1$
@@
    A multigraph $G(\Pi)$ is obtained by identifying the vertices in $\mathcal{G}$ according to the partition $\Pi$. The multigraph $G(\Pi)$ has $r$ edges and $k$ vertices, and each vertex has a degree of at least 2. Also, we focus on the ones that are loop-free, since $A$ is diagonal-free and any loop in $G(\Pi)$ gives a zero entry in the sum.
@@
$\mathbb{E}|X_{i}|^{l}\leq p_{i}(K{l})^{l}\leq p_{i}(eK)^{l}l!,$
@@
The proof of Lemma 5.2 is deferred to Section 5.1.1. It is worth noting that the bound on $\mathcal{S}_{G(\Pi)}$ in Lemma 5.2 is uniform and independent of $G(\Pi)$. To proceed from (5.5), we now need to estimate
@@
    $\sum_{\phi_{i}:[2r_{i}]\to[k_{i}]}\prod_{\tau=1}^{k}\deg_{G(\phi_{i})}(\tau)!$ $\leq\sum_{d_{1},\dots,d_{k_{i}}}\binom{2r_{i}}{d_{1},\dots,d_{k_{i}}}\cdot\prod_{\ell=1}^{k_{i}}d_{\ell}!$ (5.11)
@@
    $\leq\binom{2r_{i}}{r_{i}}(2r_{i})!$ (5.13)
    $=\left(\frac{(2r_{i})!}{r_{i}!}\right)^{2}$ (5.14)
    $\leq(2r_{i})^{2r_{i}}.$ (5.15)
Combining all the analysis of $\Pi$ above, with $r_{i}\geq 2$ we have
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ $\leq\frac{2^{r+k}}{k!}\max_{\text{component partition}}\left(\prod_{i=1}^{t}\binom{r}{r_{i}-1}\binom{k}{k_{i}}\right)\left(\prod_{i=1}^{t}2^{2r_{i}}r_{i}^{2r_{i}}\right)$
$\leq\frac{2^{r+k}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{er}{r_{i}-1}\right)^{r_{i}-1}\left(\frac{ek}{k_{i}}\right)^{k_{i}}2^{2r_{i}}r_{i}^{2r_{i}}$
$\leq\frac{2^{5r}e^{2r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r}{r_{i}}\right)^{r_{i}-1}\left(\frac{k}{k_{i}}\right)^{k_{i}}r_{i}^{2r_{i}}$
$\leq\frac{2^{5r}e^{2r}r^{2r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r}{r_{i}}\right)^{r_{i}-1}\left(\frac{k}{k_{i}}\right)^{k_{i}}\left(\frac{r_{i}}{r}\right)^{2r_{i}}$
$\underset{(\text{a})}{\leq}\frac{2^{5r}e^{2r}r^{2r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r_{i}}{r}\right)^{r_{i}-k_{i}+1}.$ (5.16)
@@
Here, the last inequality follows from Lemma A.2 and that the maximum of $\prod_{i}z_{i}^{z_{i}}$ is attained when $t-1$ of the $z_{i}$’s are equal to 1, and the remaining one equals $Z-(t-1)$. Therefore, we have
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ $\leq\frac{2^{5r}e^{2r}r^{2r}}{k!}\cdot\frac{(r-k+1)^{r-k+1}}{(r-k+t)^{r-k+t}}$ (5.18)
$\leq 2^{5r}e^{2r+k}r^{2r}k^{-k}{t^{-(t-1)}}$ (5.19)
$\leq 2^{5r}e^{3r}r^{2r-k-t+1}\left(\frac{r}{k}\right)^{k}\left(\frac{r}{t}\right)^{t}$ (5.20)
$\leq 2^{6r}e^{(3+2/e)r}r^{2r-k-t}.$ (5.21)
@@
$\leq(64e^{(5+2/e)r}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}\cdot r^{2r-k-t}$ (5.24)
$\leq(64e^{(5+2/e)r}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}(r^{2}\|A\|_{\max})^{r-k}(r\gamma_{1,\infty})^{k-2t}(\sqrt{r}\sigma)^{2t}$ (5.25)
$\leq(64e^{(5+2/e)r}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}3^{r}\max\left\{\frac{(r^{2}\|A\|_{\max})^{r}}{r/(r-k)},\frac{(r\gamma_{1,\infty})^{r}}{r/(k-2t)},\frac{(\sqrt{r}\sigma)^{r}}{r/2t}\right\}$ (5.26)
$\leq(192e^{(5+2/e)}K^{2})^{r}\cdot\frac{r^{2}}{2}\max\left\{{(r^{2}\|A\|_{\max})^{r}},{(r\gamma_{1,\infty})^{r}},{(\sqrt{r}\sigma)^{r}}\right\}$ (5.27)
$\leq(C_{1}K^{2})^{r}\max\left\{{(r^{2}\|A\|_{\max})^{r}},{(r\gamma_{1,\infty})^{r}},{(\sqrt{r}\sigma)^{r}}\right\},$ (5.28)
where we can take absolute constant $C_{1}=384e^{(5+2/e)}$. Here in the deduction we use the fact that $r^{2}\leq 2^{r}$ if $r\in\{2\}\cup[4,+\infty)$.
5.1.1 Proof of Lemma 5.2
@@
$\mathcal{S}_{G}:=\sum_{s\in[n]^{V(G)}}\left(\prod_{\{\pi_{1},\pi_{2}\}\in E(G)}a_{s_{\pi_{1}}s_{\pi_{2}}}\cdot\prod_{\pi\in V(G)}p_{s_{\pi}}\right)$ (5.29)
@@
    For each deleted edge, say $\{u,v\}$, we bound the corresponding $a_{uv}\leq\|A\|_{\max}$, and we will delete $r-k$ many edges this way in total, which implies that we delete a total of $r-k$ edges from $G$. These deleted edges collectively contribute a factor of $\|A\|_{\max}^{r-k}$ to bound $\mathcal{S}_{G}$, i.e.,
@@
5.2 Proof of Lemma 5.1 for any $\alpha>0$
We now present the essential modifications to the case $\alpha=1$ required to extend the proof of Lemma 5.1 to arbitrary $\alpha\in(0,\infty]$, highlighting the differences.
First of all, the moments of sparse $\alpha$-subexponential random variables satisfy
$\mathbb{E}|X_{i}|^{l}\leq p_{i}K^{l}l^{\frac{l}{\alpha}}.$
For $\alpha\geq 1$, we simply bound $\mathbb{E}|X_{i}|^{l}\leq p_{i}K^{l}l^{l}$, which leads to the same moment bounds as in the case where $\alpha=1$:
$\|X^{T}AX\|_{L_{r}}\leq C_{1}K^{2}\max\left\{r^{2}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}.$ (5.37)
Next, we will focus on the case where $\alpha\in(0,1)$. From the moment bounds
$\mathbb{E}|X_{i}|^{l}\leq p_{i}K^{l}l^{\frac{l}{\alpha}}\leq p_{i}(eK)^{l}l!\cdot l^{(\frac{1}{\alpha}-1)l},$
@@
$\mathbb{E}|X^{T}AX|^{r}$ $\leq(64e^{5+2/e}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}r^{2r-k-t}\cdot\big{(}2^{2r}r^{2(r-k+1)}\big{)}^{\frac{1}{\alpha}-1}$ (5.40)
$\leq(2^{2(\frac{1}{\alpha}-1)}\cdot 64e^{(5+2/e)}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}r^{2r-k-t}\cdot r^{2(\frac{1}{\alpha}-1)(r-k+1)}$ (5.41)
$\leq(2^{\frac{2}{\alpha}}\cdot 16e^{5+2/e}K^{2})^{r}r^{2(\frac{1}{\alpha}-1)}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}\cdot r^{\frac{2}{\alpha}r-(\frac{2}{\alpha}-1)k-t}$ (5.42)
$=(2^{\frac{2}{\alpha}}\cdot 16e^{5+2/e}K^{2})^{r}r^{2(\frac{1}{\alpha}-1)}\sum_{k=2}^{r}\sum_{t=1}^{k/2}(r^{\frac{2}{\alpha}}\|A\|_{\max})^{r-k}(r\gamma_{1,\infty})^{k-2t}(\sqrt{r}\sigma)^{2t}$ (5.43)
$\leq(2^{\frac{2}{\alpha}}\cdot 16e^{5+2/e}K^{2})^{r}r^{2(\frac{1}{\alpha}-1)}\cdot\frac{r^{2}}{2}\cdot 3^{r}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}^{r}$ (5.44)
$\leq(2^{\frac{3}{\alpha}}\cdot 48e^{5+2/e}K^{2})^{r}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}^{r}.$ (5.45)
Here we again use Young’s inequality for the products of three terms.
Hence, we have
@@
for any positive even integer $r$ by choosing $C_{\alpha}=48e^{5+2/e}\cdot 8^{1/\alpha}$ (in the case $\alpha\in(0,1]$). Combining this bound with $C_{1}=384e^{(5+2/e)}$ in the case of (5.37), by taking $C_{\alpha}=48e^{5+2/e}\cdot 8^{\max\{1,1/\alpha\}}$ in general completes the proof of Lemma 5.1.
5.3 Proof of Theorem 1
$\leq\left(\frac{1}{t}C_{\alpha}K^{2}\max\left\{r^{\max\{2,\frac{2}{\alpha}\}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}\right)^{r}.$ (5.47)
For any $t>0$, if
$\min\left\{\frac{t^{2}}{(C_{\alpha}e)^{2}K^{4}\sigma^{2}},\frac{t}{C_{\alpha}eK^{2}\gamma_{1,\infty}},\left(\frac{t}{C_{\alpha}eK^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}<2,$ (5.48)
then in the inequality in Theorem 1, there is
$\exp\left(-\min\left\{\frac{t^{2}}{(C_{\alpha}e)^{2}K^{4}\sigma^{2}},\frac{t}{C_{\alpha}eK^{2}\gamma_{1,\infty}},\left(\frac{t}{C_{\alpha}eK^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}\right)\geq e^{-2}.$
The conclusion of Theorem 1 holds since the right-hand side of the bound is at least 1.
$2\leq r\leq\min\left\{\frac{t^{2}}{(C_{\alpha}e)^{2}K^{4}\sigma^{2}},\frac{t}{C_{\alpha}eK^{2}\gamma_{1,\infty}},\left(\frac{t}{C_{\alpha}eK^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\},$
we have that
$\frac{1}{t}C_{\alpha}K^{2}\max\left\{r^{\max\{2,\frac{2}{\alpha}\}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}\leq\frac{1}{e},$
and further
$\mathbb{P}\left(\left|\sum_{i\not=j}a_{ij}X_{i}X_{j}\right|\geq t\right)$
$\leq e^{-r}$
$\leq\exp\left(-\left(\min\left\{\frac{t^{2}}{(C_{\alpha}e)^{2}K^{4}\sigma^{2}},\frac{t}{C_{\alpha}eK^{2}\gamma_{1,\infty}},\left(\frac{t}{C_{\alpha}eK^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}-2\right)\right)$
$=e^{2}\exp\left(-c_{\alpha}\min\left\{\frac{t^{2}}{K^{4}\sigma^{2}},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}\right),$
by choosing $c_{\alpha}=\min\{{(C_{\alpha}e)^{-2}},{(C_{\alpha}e)^{-1}},{(C_{\alpha}e)^{-\min\{\frac{\alpha}{2},\frac{1}{2}\}}}\}=(C_{\alpha}e)^{-2}$. The proof of Theorem 1 is complete.
5.4 Proof of Theorem 2
Theorem 2 is established through the following moment estimates of the linear form, which we shall prove first.